OpenStudy (anonymous):

1 + cos(18x) = _____.

zepdrix (zepdrix):

I'm not really sure what they're looking for here... I guess you could apply the Cosine Double Angle Formula,\[\Large\rm =1+\left[\color{orangered}{\cos(18x)}\right]\]\[\Large\rm =1+\left[\color{orangered}{2cos(9x)-1}\right]\]Which simplifies a lil bit,\[\Large\rm =2\cos(9x)\]

zepdrix (zepdrix):

Woops that should be a square on the cosine >.< my bad.

zepdrix (zepdrix):

\[\Large\rm =1+\left[\color{orangered}{2\cos^2(9x)-1}\right]\]

zepdrix (zepdrix):

\[\Large\rm =2\cos^2(9x)\]

zepdrix (zepdrix):

Is that what they wanted or no? :U

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yes it is

zepdrix (zepdrix):

cool c: